OpenStudy (anonymous):

simplify 4x-20 over 4x

OpenStudy (pfenn1):

\[4x-20 \over 4x\] factor the 4 out of the numerator.

OpenStudy (pfenn1):

No it doesn't.

OpenStudy (pfenn1):

\[\frac{4x-20}{4x}=\frac{4(x-5)}{4x}\]

OpenStudy (anonymous):

your right.. i messed up